OpenStudy (anonymous):

solve the equation 4x+1/8x-1=3/5

OpenStudy (mertsj):

\[\frac{4x+1}{8x-1}=\frac{3}{5}\]

OpenStudy (mertsj):

Cross multiply

OpenStudy (anonymous):

ok so 20x+5=24x-3?